Find an equation of the line having the given slope and containing the given point. m=4/5, (9,-8)
Answer by Cromlix(4381)

y - (-8) = 4/5 (x - 9)
y + 8 = 4/5 (x - 9)
5y + 40 = 4x - 36
5y = 4x - 76 or y = 4/5x - 76/5